In a survey of a group of 105 movie-goers, 60 liked the movie Aquaman, 30 liked the movie Roma and 20 liked both movies. How many of the surveyed did not like either of the movies? 40 35 0 10 70
Answer #1

Let the event A denotes for the movie-goers like Aquaman

The event B denotes for the movie-goers like Roma.

N denotes total number of movie-goers.

Here, N = 105

n(A) = 60

n(B) = 30

n(A and B) = 20

n(A or B) = n(A) + n(B) - n(A and B)

= 60 + 30 - 20

= 70

And number of surveyed did not like either of the movies

= 105 - 70

= 35